OpenStudy (anonymous):

Evaluate m-2n/6 if m=-20 and n=5 Would it be: -18 -5 or -15 ?

jhonyy9 (jhonyy9):

will be (-20-2*5)/6 = (-20-10)/6 = -30/6 = -5 hope you understand it sure easy